find the perimeter of a rectangle whose length is 5cm less than twice its width.
Answer by mananth(16949)

let width =x
twice width = 2x
5 less than twice width = 2x-5
length = 2x-5
Perimeter = 2*(L+W)
=2(2x-5+x)
=2(3x-5)